OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (4, –2) and has a slope of –3? x + 3y = 10 x + 3y = –10 3x + y = 10 3x + y = –10

OpenStudy (lgbasallote):

use point slope form \[y - y_1 = m(x-x_1)\] your m here is -3 because it's the slope x1 is 4 because it's the x coordinate y1 is -2 because it's the y coordinate \[y - (-2) = -3(x-4)\] \[y+2 = -3x + 12\] \[y = -3x + 12 - 2\] \[y = 3x +10\]

OpenStudy (anonymous):